OpenStudy (gokuporter):

Write an equation in slope intercept form for the line passing through (5,1) and (2,10)

OpenStudy (anonymous):

is there a picture?

OpenStudy (gokuporter):

no?

OpenStudy (gokuporter):

these are the options A) y=3x-16 B)y=-3x-16 C)y=-3x+16 D)y=3x+16

OpenStudy (anonymous):

Mind if I help?

OpenStudy (gokuporter):

yes

OpenStudy (anonymous):

ok well the slope is -3 meaning that it is either B or C, and if you put the two points on the line, I believe that +16 would be the y-intercept, therefore I think it's C

OpenStudy (gokuporter):

how did you get that

OpenStudy (anonymous):

Do you mean how did i get the slope?

OpenStudy (gokuporter):

no wait i know how to find the slope now

OpenStudy (anonymous):

ok, so are you good?

OpenStudy (gokuporter):

i got a different answer

OpenStudy (anonymous):

To find slope do (y2-y1)/x2-x1

OpenStudy (gokuporter):

i got -2 for slope

OpenStudy (anonymous):

show me your steps

OpenStudy (anonymous):

i'm going to draw what I did and show you how I got -3

OpenStudy (gokuporter):